Determine the missing binomial in the following: (√3 – 2) ( ) = -1
Answer by dolly(163)

Let the missing binomial be x
Then given (√3 – 2)(x) = -1
dividing both the sides by (√3 – 2),
==> x = -1/(√3 – 2)
Multiply and divide the right side fraction by the conjugate of the denominator
which is (√3 + 2)
==> x = -1((√3 + 2) / (√3 + 2)(√3 – 2)
==> x = -1(√3 + 2)/ (3 - 4) [applying (a+b)(a-b) = a^2 - b^2]
==> x = -1(√3 + 2) / -1
==> x = (√3 + 2)
Thus the missing binomial in the given equation is (√3 + 2)